Fermi National Laboratory
 

Accelerator Update for January 22, 2007 - January 24, 2007

 

Monday January 22

The day shift began with the Tevatron (TeV) being prepared for access, with Operations monitoring a Recycler stash (97.3E10), with the Antiproton Source (Pbar) stacking (20.9E10), with the Linac H- Source running at a lower than normal operating voltage to prevent sparking, with Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) preparing to access their collision halls.

The TeV was ready for access by 8:40 AM for maintenance work.

At 9:44 AM, NuMI tripped off for the first of two times in eight minutes. It reset both times.

At 10:28 AM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum pump (B:IPRF17) trip. Everything reset.

Operations reported at 10:54 AM that the NuMI plus slipstacking events were running every 2 seconds due to the H- Source running with a lower output.

At 11:29 AM, Operations turned off the Preaccelerator so that Linac technicians could replace the H- Source water resistor. This work halted all beam.

Operations reported at 12:02 PM that everyone was out of the TeV and that operators had started the TeV turn on.

Operations started ramping the TeV at 1:33 PM.

Operations put the TeV through a dry squeeze at 1:48 PM.

Linac techs completed the H- Source water resister replacement at 1:50 PM. Operations established beam to the Linac absorber at 1:59 PM. With tuning, the source resumed its normal output.

Operations established beam to Pbar and NuMI at 2:06 PM.

TeV experts began a proton beam study at 2:28 PM. CDF and D0 were notified of this study.

Operators and experts established slow spill through the Switchyard (SY) to Meson at 3:18 PM.

Operations reported at 3:49 PM that the Proton Source was putting out its full intensity.

At 3:49 PM, a MI RF station (MIRF10) tripped off on a 480V 3-phase absent indication. An RF expert investigated.

Meson experts reported at 4:02 PM that they were running beam to the MTest area.

Recycler experts completed a study at 4:06 PM.

TeV studies ended at 4:22 PM.

Operations loaded the TeV with a proton only wet squeeze. The beam was intentionally aborted at 4:59 PM.

Operations halted beam to Meson. SY and Meson experts had completed their studies.

Operations put the TeV into shot setup at 6:49 PM

At 7:04 PM, Operations reported a power glitch. This put the shot setup on hold. Operators reported that feeder 36 had faulted and the Duty Electrician confirmed it. The Cryo system technicians from CHL reported that one of their big helium compressors tripped off. The Meson LCW also tripped off. Operations contacted the necessary people.

CHL reported at 7:40 PM that their helium compressor was back on and that they were providing liquid helium to the TeV, but they advised Operations not to ramp the TeV for an hour.

The High Voltage Coordinator reported at 8:41 PM that feeder 36 suffered a fault. He and the Duty Electrician were inspecting transformers.

The High Voltage Coordinator reported at 9:51 PM that he found the fault at the Site 50 transformer. He isolated Site 50 and 52 from feeder 36.

The High Voltage Coordinator energized feeder 36 at 11:27 PM.

At 11:28 PM, the Cryo Coordinator reported that there was a warm bubble of gas in the TeV. It would take another 30 minutes to work its way out.

Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.

Tuesday January 23

The midnight shift began with Operations monitoring the stash (235.8E10), with Pbar stacking (54.5E10), and with NuMI taking beam.

Operations put the TeV through a dry squeeze at 12:29 AM.

Operations put the TeV into a shot setup at 1:29 AM.

Operations established store 5193 at 3:31 AM with an initial luminosity of 238.18E30.

At 4:25 AM, Operations reported another power glitch. It appeared to be feeder 36 again. The store remained stable. Operations contacted the necessary people.

The High Voltage Coordinator came in at 6:50 AM and began hunting for the new feeder fault.

At 6:57 AM, NuMI tripped off for the first of four times in 11 minutes due to power supply regulation problems on two Lambertsons and a corrector. Operators notified experts.

Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.

At 8:39 AM, CHL reported another power glitch. This glitch tripped off their nitrogen plant.

The High Voltage Coordinator confirmed a third feeder 36 fault at 8:45 AM. He was still investigating.

Operations established slow spill to Meson 9:40 AM.

The Duty Electrician reported at 10:42 AM that they had isolated the faults and were energizing feeder 36.

Operations rebooted the MI low-level RF system at 12:49 PM.

A power supply expert completed the NuMI work at 12:54 PM.

A MI expert began an hour and a half slipstacking study at 1:07 PM.

Operations began sending slow spill beam to Meson at 1:31 PM.

Operations reported at 4 PM that the Recycler had stashed three times during the day shift.

The SY/Meson experts reported that they were finished using the beam at 5:02 PM.

A Pbar expert began a study at 5:26 PM.

At 6:44 PM, Operations reported that the Pbar lithium lens located at AP0 was tripping on an external interlock. This happened three more times over the next 45 minutes. It reset each time.

TeV experts began a Tevatron Electron Lens (TEL2) study at 10 PM.

At 11:48 PM, the MI-10 line suffered a vacuum burst. The vacuum recovered, but then at 11:52 PM the MI-8 vacuum began to cascade. Operations halted MI beam so the vacuum could recover. MI beam resumed at 11:59 PM.

Operations reported at midnight that the Recycler had stashed two times during the evening shift.

Wednesday January 24

The midnight shift began with Operations monitoring store 5193 (41.19E30) and the stash (329.5E10), with Pbar stacking (17.4E10), and with NuMI taking beam.

At 1:47 AM, KRF1 tripped off. It reset okay.

At 1:59 AM, KRF5 tripped off. It reset.

Operations terminated store 5193 at 4 AM.

Operations put the TeV into shot setup at 4:14 AM.

Operations established store 5194 at 5:57 AM with an initial luminosity of 273.17E30. This was the second highest store for luminosity.

Operations reported at 7:58 AM that the Recycler had stashed two times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The plans for Thursday are to allow D0 to have a ten-hour access to repair its detector. The Run Coordinator will allow maintenance work in all machines except for MI and Recycler due to the stash.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/24/2007    email Fermilab